# GRAVITY for MATISSE

### v1

@@ 2.6 Visibility observations
An absolute flux measurement is required to convert coherent flux measurements to visibilities. Owing to the stronger thermal background in the mid-infrared, this measurement tends to be more difficult to obtain accurately; this has been a challenge for all thermal interferometers. The Keck Interferometer Nuller (Serabyn et al., 2012) could not rely on chopping secondary mirrors on both Keck telescopes and had to implement an interferometric chopping strategy on the short baselines corresponding to the two halves of a same telescope. The MIDI+FSU experiment (Müller et al., 2014) focused on faint coherent flux measurements and did not attempt to measure chopped photometry. Large Binocular Telescope Interferometer (LBTI), where PhaseCam (Defrère et al., 2014) is the HK band fringe tracker and NOMIC (Hoffmann et al., 2014) / LMIRcam (Skrutskie et al., 2010) are the thermal science cameras, is probably the facility closest to GRA4MAT.
@@
The situation could only be improved after the upgrade of the fringe tracker real-time controller (Abuter et al., 2016; Nowak et al., 2024) in November 2022. In January 2023, an active chopping state was integrated into the new real-time software. The chopping of the fringe tracker behaves like any other chopping element of VLTI. Its sequence is fully specified by an absolute start time, a period, and a sky-to-target ratio. When chopping is active, the timing of the group delay control loop set point is adjusted to fall right at the end of each TARGET phase, just before the transition to SKY. This was chosen because this is when the fringe tracker has the highest probability of having a correct estimation of the dispersion between phase delay and group delay. Updating the set point at the start of the TARGET phase would have required to first wait for the dispersion estimators to build SNR, increasing further the time needed to recover proper fringe tracking performance. For this scheduling strategy to work, the dispersion must remain within for the duration of a full chopping cycle.
Whereas the target-to-sky transition happens at the same time for all chopping participants, the sky-to-target transition must happen in sequence. First, the chopping mirror (deformable mirror on ATs and telescope secondary mirror on UTs) must move back to the target position, then the adaptive optics must close the loop and take a few iterations to deliver the correction level needed for the fringe tracker to resume control. For the current implementation, this delay is set to after measuring the time it takes for the flux injected into the fringe tracker to reach back its typical level. The fringe tracker finally needs some additional time to reacquire the fringe identified before the transition to sky. In total, we implemented a + delay from the start of the target phase to when the L band starts its integration. The majority of the delay is consumed by the fringe tracker slewing back to the reference fringe. This constraint is set by the length of the fringe tracker group delay averaging window of 150 samples or at a detector period of and by the speed limit of the group delay corrections. The maximum chopping frequency is set to ¡ (¿ on sky and ¿ on target) and practically limited by the fringe recovery time. The fixed + delay has been tuned for the period of the fringe tracker. As such, chopping with GRA4MAT is not supported when the fringe tracker operates slower, and the following limits must be respected: K ¡ with the ATs, and K ¡ with the UTs.
@@ 2.7 Data collection
The GRAVITY fringe tracker supports the MATISSE performance at two levels. First, it actively stabilizes the optical path difference and improves the MATISSE coherence beyond the atmospheric limits, up to the background limited integration times of MATISSE. Second, it contributes chromatic complex coherence and flux in K band to the data recorded by MATISSE. This information is generated directly by the fringe tracker real-time controller, published through the Reflective Memory Network (RMN), and recorded by the RMN recorder (Abuter et al., 2008) under the control of the MATISSE parallel exposures. These records are part of the dataset sent to the ESO archive and used by the instrument pipeline for coherent integrations and frame selection.
@@ 2.8 Data reduction
The pipeline first uses the GRAVITY telemetry to detect fringe jumps, i.e. when the fringe tracker group delay moves by more than half a wavelength. The recorded MATISSE frames that contain jumps are flagged with TARTYP=J, and dismissed from the remaining computations of the pipeline. The extent of the flagging depends on the MATISSE spectral bands. In N band, all the frames of a complete 10-steps modulation cycle are rejected, corresponding to a period of in LOW resolution and in HIGH. In L band, only the frame corresponding to the detected jump is rejected. Note that even if fringe jumps are detected in the group delay of one baseline, all baselines are affected by the flag.
In the N band, the thermal background contribution dominates many astrophysical sources. The N-band detector integration time is then set to avoid any thermal background saturation: in LOW spectral resolution, in HIGH. As such, coherent integration in the pipeline are necessary to improve the signal-to-noise of interferometric observables. In the GRA4MAT mode, the estimation of the N-band phase is derived from the fringe tracker K-band measurements. Even if the fringes are stabilized by the fringe tracker, the N-band OPD fluctuates at a level of a few microns, due to the dispersive water vapour turbulence. The MATISSE pipeline follows the procedure of Koresko et al. (2006) to extrapolate the K-band measurements to the N-band: the differential column density of water vapour is estimated from the K-band phase delay and group delay measurements, and linearly converted into a prediction of the phase delay and group delay in the N band. We recall the equations involved in this conversion:
where the phase delay , the group delay , and the dispersion are in unit of optical length. is the water vapour refractivity and is its group equivalent. Table 2 gives numerical values for the above parameters, derived from the refractive indexes of Colavita et al. (2004). Figure 2 illustrates the K-to-N prediction process on a target bright enough in N band to estimate the phase directly from the MATISSE data.
A similar K-to-L correction is currently not implemented in the MATISSE pipeline. Since the water vapour decorrelation from K band to L band is not as strong as to N band (Table 5 in Colavita et al., 2004), such a correction was not deemed critical to the project. Recent developments have however shown that a residual phase jitter correction would certainly help stabilise the MATISSE transfer function, and open the door to higher contrast observations.
Band | ||||
K () | 0.000 | 1.000 | ||
L () | 0.256 | 1.429 | ||
M () | 0.745 | 3.312 | ||
N () | 2.780 | 7.824 |
@@ 3 Fringe tracking in the GRA4MAT context
In this section, we present aspects of the GRAVITY fringe tracker performance, when supporting MATISSE, an instrument operating at a different wavelength.
@@
The GRAVITY fringe tracker uses the main VLTI delay lines as OPD actuators when operating with GRA4MAT. The delay lines are known to be affected by extra latency, reducing the OPD control performance when compared with the internal TTP actuator of GRAVITY. Fig. 3 compares the responses of the two actuators and shows a additional latency. This actuator response, a necessary parameter of the controller implementation (see Nowak et al., 2024), is updated each time the fringe tracker switches between GRAVITY and GRA4MAT. The consequences for MATISSE are however limited, since it operates at longer wavelength, and this small fringe tracking performance degradation is by far not the dominant factor.
@@
As of 2023, GRA4MAT operates with NAOMI (Woillez et al., 2019) on the ATs and MACAO (Arsenault et al., 2004) on the UTs. We recall that NAOMI has significantly smaller sub-apertures than MACAO ( versus ). In the high flux regime, the Strehl delivered by NAOMI is therefore higher than the one delivered by MACAO. In addition, due to a reduced level of turbulence aliasing in the sub-apertures of the wavefront sensor, the NAOMI performance does not degrade as quickly as MACAO when the seeing conditions become poorer than median. Because of this, we obtain very different fringe tracking behaviour when comparing ATs and UTs. On the ATs, fed with a higher and more stable Strehl, the fringe tracker can continuously track the same fringe without jumping. Whereas on the UTs, the lower and more variable Strehl results in flux dropouts on the fringe tracker, which significantly increase the probability of fringe jumps. The connection between median Strehl, flux losses and fringe jumps is best illustrated in Woillez et al. (2019).
@@
The fringe jumps were not as bad for GRAVITY standalone as they are for GRA4MAT. GRAVITY operates a spectrometer in the same band as its fringe tracker, and a fringe jump only has limited effects on the sides of the K band. As shown in Fig. 4, the situation is very different for GRA4MAT where a fringe jump in K band generates a fringe phase shift very different from : at (L band), at (M band), and at (N band). Also, since the fringe jumps can happen anywhere inside a DIT, the resulting fringe contrast becomes highly variable and impossible to calibrate without knowledge of when these jumps happen.
@@
Initially, the GRAVITY fringe tracker did not have any mechanism in place to mitigate the impact of the fringe jumps. As described in Lacour et al. (2019), the group delay controller had a dead zone of (or ), leaving room for 3 stable tracking solutions that fringe jumps could randomly explore. In November 2022, the real-time controller was upgraded, replacing the slower VME-based implementation (Lacour et al., 2019) by a faster Linux workstation (Abuter et al., 2016; Nowak et al., 2024). This faster fringe tracker made possible the deployment of an improved group delay control algorithm that reduces the dead zone to less than a wavelength, leaving room for exactly one stable tracking solution for the duration of a GRAVITY spectrometer or MATISSE L-band DIT. The constant set point of the group delay controller is now updated in between the MATISSE DITs. The time between updates must remain below the evolution timescale of the dispersion caused by dry air and wet air turbulence. If not, when the dispersion shifts by more than a wavelength, the fringe tracker jumps without correction.
@@
With this new implementation, the fringe jumps happen as often as before, but are detected and corrected in less than half the group delay estimator filter length of 150 elements, i.e. when the integration period is the standard . Figure 4 illustrates the theoretical improvement of the fringe contrast, and fringe contrast stability, achieved with the new implementation.
@@
The amount of water vapour turbulence varies with time and makes the maximum L-band DIT dependent on the atmospheric conditions. We however do not have yet any analysis comparing the dispersion fluctuations to the precipitable water vapour or humidity, that would let the observer identify when conditions are compatible with absolutely no dispersion-induced fringe jumps. Preliminary observations indicate that a line-of-sight precipitable water vapour below could be an acceptable threshold.
@@
We conclude this section by recalling that the K-band limiting magnitudes of GRA4MAT are identical to GRAVITY operating in one of the off-axis modes. The expected 3 magnitudes difference between ATs and UTs remains unobserved for now; this is a consequence of the lower relative performance of the UT adaptive optics.
@@ 4 MATISSE performance improvements
In this section, we present aspects of the MATISSE performance improvements obtained when operating with the GRAVITY fringe tracker.
4.1 Full LM-bands spectral coverage
MATISSE employs a HAWAII-2RG detector for the LM-bands observations. Its readout speed is limited by a slow pixel clock of . This corresponds to a full detector readout time of , well beyond the atmosphere coherence time. Windowing the detector is the only solution to reduce the readout time, at the expense of spectral resolution and/or wavelength coverage. At an integration time of , within the L-band atmospheric coherence time, a full LM-bands spectral coverage can only be achieved in LOW spectral resolution. For higher spectral resolutions (MEDIUM and HIGH), only a small fraction of the wavelength coverage can be obtained. The highest spectral resolution (HIGH+) is not even offered with MATISSE standalone.
The situation changes completely with GRA4MAT. When fringes are stabilized, the LM-bands detector integration time can be increased to without coherence loss and the full spectrum can be read. Illustrated in Figure 5, the observation of the spectroscopically-rich binary Be star Cen showcases this improved wavelength coverage. To obtain the same spectral coverage without a fringe tracker, one would need a sequence of about 11 consecutive observations. This represents a gain in observing efficiency.
4.2 Improved LM-bands sensitivity
When the MATISSE fringes are stabilized by the GRAVITY fringe tracker, one would ideally want to increase the DITs to a regime where the detector readout noise becomes negligible with respect to the thermal background noise. However, to keep the probability of fringe jumps to a minimum within each DIT, and to preserve the OPD modulation capability within an exposure, having the thermal background noise slightly above the detector readout noise is sufficient. Table 3 gives the DITs at various wavelengths that correspond to the background noise matching the readout noise.
L-band | LOW | MEDIUM | HIGH | HIGH+ |
@@ 4.2 Improved LM-bands sensitivity
In Fig. 6, we show a model of the error on the chromatic differential phase in L band, as a function of the target flux, for each spectral resolution. This model is identical to the one used in Lopez et al. (2022, see Fig. 11 therein), but with DITs extended to the GRA4MAT regime. Its validity was verified against on-sky GRA4MAT observations. In LOW spectral resolution, an increase of the detector exposure time from to , where the background noise dominates, yields a sensitivity limit improvement of . Following the ESO Call for Proposal convention adopted for MATISSE that considers chromatic phase errors of per spectral channel in exposure, we reach with the ATs and with the UTs. In MEDIUM spectral resolution, the detector exposure time can be increased further, to a background-dominated regime of , corresponding to a sensitivity limit improvement of . At higher spectral resolution, we remain with a detector exposure time of , which is around the optimum in HIGH, but well below in HIGH+. Similar levels of improvement are obtained for the L-band closure phase and for the M-band differential phase and closure phase, but not illustrated herein. We recall that the instrument is perfectly capable of observing targets fainter than the flux levels given above, when the intended science does not require the or high precision. At or , GRA4MAT reaches in LOW spectral resolution with the UTs, as illustrated in Fig. 6.
As already known since Lopez et al. (2022), at increasing flux levels the errors on the MATISSE observables decrease and reach a minimum plateau. The behaviour remains unchanged when MATISSE operates with a fringe tracker, as shown in the same Fig. 6. This observation does not help understand where the effects come from, but it probably rules out the role of residual piston. This plateau is not investigated further in this work.

### v2

@@ 2.6 Visibility observations
An absolute flux measurement is required to convert coherent flux measurements to visibilities. Owing to the stronger thermal background in the mid-infrared, this measurement tends to be more difficult to obtain accurately; this has been a challenge for all thermal interferometers. The Keck Interferometer Nuller (Serabyn et al., 2012) could not rely on chopping secondary mirrors on both Keck telescopes, and had to implement an interferometric chopping strategy on the short baselines corresponding to the two halves of a same telescope. The MIDI+FSU experiment (Müller et al., 2014) focused on faint coherent flux measurements and did not attempt to measure chopped photometry. LBTI, where PhaseCam (Defrère et al., 2014) is the HK-band fringe tracker and NOMIC (Hoffmann et al., 2014) and LMIRcam (Skrutskie et al., 2010) are the thermal science cameras, is probably the facility closest to GRA4MAT.
@@
The situation could only be improved after the upgrade of the fringe tracker real-time controller (Abuter et al., 2016; Nowak et al., 2024) in November 2022. In January 2023 an active chopping state was integrated into the new real-time software. The chopping of the fringe tracker behaves like any other chopping element of VLTI. Its sequence is fully specified by an absolute start time, a period, and a sky-to-target ratio. When chopping is active, the timing of the group delay control loop set point is adjusted to fall right at the end of each TARGET phase, just before the transition to SKY. This was chosen because this is when the fringe tracker has the highest probability of having a correct estimation of the dispersion between phase delay and group delay. Updating the set point at the start of the TARGET phase would have required to first wait for the dispersion estimators to build S/N, increasing further the time needed to recover proper fringe tracking performance. For this scheduling strategy to work, the dispersion must remain within for the duration of a full chopping cycle.
Although the target-to-sky transition happens at the same time for all chopping participants, the sky-to-target transition must happen in sequence. First the chopping mirror (deformable mirror on ATs and telescope secondary mirror on UTs) must move back to the target position, then the adaptive optics must close the loop and take a few iterations to deliver the correction level needed for the fringe tracker to resume control. For the current implementation, this delay is set to after measuring the time it takes for the flux injected into the fringe tracker to return to its typical level. The fringe tracker finally needs some additional time to reacquire the fringe identified before the transition to sky. In total, we implemented a + delay from the start of the target phase to when the L band starts its integration. The majority of the delay is consumed by the fringe tracker slewing back to the reference fringe. This constraint is set by the length of the fringe tracker group delay averaging window of 150 samples or at a detector period of and by the speed limit of the group delay corrections. The maximum chopping frequency is set to ¡ (¿ on sky and ¿ on target) and practically limited by the fringe recovery time. The fixed + delay has been tuned for the period of the fringe tracker. As such, chopping with GRA4MAT is not supported when the fringe tracker operates more slowly, and the following limits must be respected: K ¡ with the ATs, and K ¡ with the UTs.
@@ 2.7 Data collection
The GRAVITY fringe tracker supports the MATISSE performance at two levels. First, it actively stabilises the optical path difference and improves the MATISSE coherence beyond the atmospheric limits up to the background limited integration times of MATISSE. Second, it contributes chromatic complex coherence and flux in the K band to the data recorded by MATISSE. This information is generated directly by the fringe tracker real-time controller, published through the reflective memory network (RMN), and recorded by the RMN recorder (Abuter et al., 2008) under the control of the MATISSE parallel exposures. These records are part of the dataset sent to the ESO archive and used by the instrument pipeline for coherent integrations and frame selection.
@@ 2.8 Data reduction
The pipeline first uses the GRAVITY telemetry to detect fringe jumps (i.e. when the fringe tracker group delay moves by more than half a wavelength). The recorded MATISSE frames that contain jumps are flagged with TARTYP=J, and dismissed from the remaining computations of the pipeline. The extent of the flagging depends on the MATISSE spectral bands. In N band, all the frames of a complete ten-step modulation cycle are rejected, corresponding to a period of in low resolution and in high. In L band, only the frame corresponding to the detected jump is rejected. We note that even if fringe jumps are detected in the group delay of one baseline, all baselines are affected by the flag.
In the N band, the thermal background contribution dominates many astrophysical sources. The N-band detector integration time is then set to avoid any thermal background saturation: in low spectral resolution, in high. As such, coherent integration in the pipeline are necessary to improve the S/N of interferometric observables. In the GRA4MAT mode, the estimation of the N-band phase is derived from the fringe tracker K-band measurements. Even if the fringes are stabilised by the fringe tracker, the N-band OPD fluctuates at a level of a few microns, due to the dispersive water vapour turbulence. The MATISSE pipeline follows the procedure of Koresko et al. (2006) to extrapolate the K-band measurements to the N band: the differential column density of water vapour is estimated from the K-band phase delay and group delay measurements, and linearly converted into a prediction of the phase delay and group delay in the N band. We recall the equations involved in this conversion:
Here the phase delay , the group delay , and the dispersion are in units of optical length; is the water vapour refractivity; and is its group equivalent. Table 2 gives numerical values for the above parameters, derived from the refractive indexes of Colavita et al. (2004). Figure 2 illustrates the K-to-N prediction process on a target bright enough in N band to estimate the phase directly from the MATISSE data.
A similar K-to-L correction is currently not implemented in the MATISSE pipeline. Since the water vapour decorrelation from K band to L band is not as strong as to N band (Table 5 in Colavita et al., 2004), such a correction was not deemed critical to the project. However, recent developments have shown that a residual phase jitter correction would certainly help stabilise the MATISSE transfer function, and open the door to higher contrast observations.
@@ 3 Fringe tracking in the GRA4MAT context
In this section we present aspects of the GRAVITY fringe tracker performance, when supporting MATISSE, an instrument operating at a different wavelength. The GRAVITY fringe tracker uses the main VLTI delay lines as OPD actuators when operating with GRA4MAT. The delay lines are known to be affected by extra latency, reducing the OPD control performance when compared with the internal TTP actuator of GRAVITY. Figure 3 compares the responses of the two actuators and shows a additional latency. This actuator response, a necessary parameter of the controller implementation (see Nowak et al., 2024), is updated each time the fringe tracker switches between GRAVITY and GRA4MAT. The consequences for MATISSE are limited, however, since it operates at longer wavelength, and this small fringe tracking performance degradation is by far not the dominant factor.
@@
As of 2023, GRA4MAT operates with NAOMI (Woillez et al., 2019) on the ATs and MACAO (Arsenault et al., 2004) on the UTs. We recall that NAOMI has significantly smaller sub-apertures than MACAO ( versus ). In the high flux regime, the Strehl delivered by NAOMI is therefore higher than that delivered by MACAO. In addition, due to a reduced level of turbulence aliasing in the sub-apertures of the wavefront sensor, the NAOMI performance does not degrade as quickly as MACAO when the seeing conditions become poorer than median. Because of this, we obtain very different fringe tracking behaviour when comparing ATs and UTs. On the ATs, fed with a higher and more stable Strehl, the fringe tracker can continuously track the same fringe without jumping, whereas on the UTs the lower and more variable Strehl results in flux dropouts on the fringe tracker, which significantly increase the probability of fringe jumps. The connection between median Strehl, flux losses, and fringe jumps is best illustrated in Woillez et al. (2019).
@@
The fringe jumps were not as bad for GRAVITY standalone as they are for GRA4MAT. GRAVITY operates a spectrometer in the same band as its fringe tracker, and a fringe jump only has limited effects on the sides of the K band. As shown in Fig. 4, the situation is very different for GRA4MAT where a fringe jump in K band generates a fringe phase shift very different from : at (L band), at (M band), and at (N band). In addition, since the fringe jumps can happen anywhere inside a DIT, the resulting fringe contrast becomes highly variable and impossible to calibrate without knowing when these jumps happen.
@@
Initially, the GRAVITY fringe tracker did not have any mechanism in place to mitigate the impact of the fringe jumps. As described in Lacour et al. (2019), the group delay controller had a dead zone of (or ), leaving room for three stable tracking solutions that fringe jumps could randomly explore. In November 2022, the real-time controller was upgraded, replacing the slower VME-based implementation (Lacour et al., 2019) by a faster Linux workstation (Abuter et al., 2016; Nowak et al., 2024). This faster fringe tracker made possible the deployment of an improved group delay control algorithm that reduces the dead zone to less than a wavelength, leaving room for exactly one stable tracking solution for the duration of a GRAVITY spectrometer or MATISSE L-band DIT. The constant set point of the group delay controller is now updated between the MATISSE DITs. The time between updates must remain below the evolution timescale of the dispersion caused by dry air and wet air turbulence. If not, when the dispersion shifts by more than a wavelength, the fringe tracker jumps without correction.
@@
With this new implementation, the fringe jumps happen as often as before, but are detected and corrected in less than half the group delay estimator filter length of 150 elements (i.e. when the integration period is the standard ). Figure 4 illustrates the theoretical improvement of the fringe contrast, and fringe contrast stability, achieved with the new implementation.
@@
The amount of water vapour turbulence varies with time and makes the maximum L-band DIT dependent on the atmospheric conditions. However, we do not yet have any analysis comparing the dispersion fluctuations to the precipitable water vapour or humidity, which would let the observer identify when conditions are compatible with absolutely no dispersion-induced fringe jumps. Preliminary observations indicate that a line-of-sight precipitable water vapour below could be an acceptable threshold.
@@
We conclude this section by recalling that the K-band limiting magnitudes of GRA4MAT are identical to GRAVITY operating in one of the off-axis modes. The expected three-magnitude difference between ATs and UTs remains unobserved for now. This is a consequence of the lower relative performance of the UT adaptive optics.
@@ 4 MATISSE performance improvements
In this section, we present aspects of the MATISSE performance improvements obtained when operating with the GRAVITY fringe tracker. A comparison between MATISSE standalone and with GRA4MAT is also presented in Table 4.
4.1 Full LM-band spectral coverage
MATISSE employs a HAWAII-2RG detector for the LM-band observations. Its readout speed is limited by a slow pixel clock of . This corresponds to a full detector readout time of , well beyond the atmosphere coherence time. Windowing the detector is the only solution to reduce the readout time, at the expense of spectral resolution and/or wavelength coverage. At an integration time of , within the L-band atmospheric coherence time, a full LM-band spectral coverage can only be achieved at low spectral resolution. For higher spectral resolutions (medium and high), only a small fraction of the wavelength coverage can be obtained. The highest spectral resolution (high+) is not even offered with MATISSE standalone.
The situation changes completely with GRA4MAT. When fringes are stabilised, the LM-band detector integration time can be increased to without coherence loss and the full spectrum can be read. Illustrated in Figure 5, the observation of the spectroscopically rich binary Be star Cen shows this improved wavelength coverage. To obtain the same spectral coverage without a fringe tracker, one would need a sequence of about 11 consecutive observations. This represents a ten-fold gain in observing efficiency.
4.2 Improved LM-band sensitivity
When the MATISSE fringes are stabilised by the GRAVITY fringe tracker, one would ideally want to increase the DITs to a regime where the detector readout noise becomes negligible with respect to the thermal background noise. However, to keep the probability of fringe jumps to a minimum within each DIT, and to preserve the OPD modulation capability within an exposure, having the thermal background noise slightly above the detector readout noise is sufficient. Table 3 gives the DITs at various wavelengths that correspond to the background noise matching the readout noise.
L band | low | medium | high | high+ |
@@ 4.2 Improved LM-band sensitivity
In Fig. 6 we show a model of the error on the chromatic differential phase in L band as a function of the target flux, for each spectral resolution. This model is identical to the one used in Lopez et al. (2022, see Fig. 11 therein), but with DITs extended to the GRA4MAT regime. Its validity was verified against on-sky GRA4MAT observations. In low spectral resolution, an increase in the detector exposure time from to , where the background noise dominates, yields a two-fold sensitivity limit improvement. Following the ESO Call for Proposal convention adopted for MATISSE that considers chromatic phase errors of per spectral channel in exposure, we reach with the ATs and with the UTs. In medium spectral resolution, the detector exposure time can be increased further, to a background-dominated regime of , corresponding to a six-fold sensitivity limit improvement. At higher spectral resolution we remain with a detector exposure time of , which is around the optimum in high, but well below in high+. Similar levels of improvement are obtained for the L-band closure phase and for the M-band differential phase and closure phase, but not illustrated here. We recall that the instrument is perfectly capable of observing targets fainter than the flux levels given above, when the intended science does not require the or high precision. At or , GRA4MAT reaches in low spectral resolution with the UTs, as illustrated in Fig. 6.
As already known since Lopez et al. (2022), at increasing flux levels the errors on the MATISSE observables decrease and reach a minimum plateau. The behaviour remains unchanged when MATISSE operates with a fringe tracker, as shown in Fig. 6. This observation does not help us to understand where the effects come from, but it probably rules out the role of residual piston. This plateau is not investigated further in this work.